# Robust Distributed Estimation: Extending Gossip Algorithms to Ranking and Trimmed Means

### v6

@@ Appendix I Further Discussion and Future Work
Ranking.
An interesting direction is to explore whether the convergence rates of GoRank are theoretically optimal, inspired by the foundational work of Boyd et al. [3]. Another important yet more challenging problem is to derive convergence rates for Asynchronous GoRank and Baseline++ under arbitrary communication graphs. Beyond convergence guarantees, it would be interesting to extend our framework to higher-dimensional ranking problems, such as estimating data depths [27]. Finally, although GoRank is currently limited to scenarios without ties, extending it to handle tied values is a natural and likely straightforward next step [23].
Robust Mean Estimation.
GoTrim opens several promising directions for future research. While currently limited to one-dimensional data, a natural extension is to multivariate settings, such as coordinate-wise trimmed means [28, 37]. The framework also generalizes naturally to robust mean computations of the form $(1/n)\sum_{i=1}^{n}f_{i}(S_{i})$, where each $f_{i}:\mathcal{S}\to\mathbb{R}$ is a local function over an arbitrary input space $\mathcal{S}$. Another compelling direction is the estimation of general $L$-statistics [33], for which GoTrim is particularly well-suited: it requires only an adaptation of the weight function $w_{n,\alpha}(\cdot)$. However, computing certain $L$-statistics, such as the median or Winsorized mean [33], remains challenging due to the need for high-accuracy rank estimation. While our current approach employs uniform sampling, more adaptive strategies—similar to Metropolis–Hastings random walks [36, 15]—could be explored to reduce the influence of corrupted nodes and improve robustness.
@@
The proposed methods offer a promising foundation for robust decentralized optimization. In particular, GoTrim could be integrated into existing mean-based optimization algorithms to enhance robustness against outliers [24, 9, 35]. Realizing this integration, however, will require further algorithmic and theoretical development. Additionally, our ranking algorithms may prove valuable in extreme value theory, where identifying rare or extreme observations is critical—especially in high-stakes domains such as finance, insurance, and environmental science [21].

### v7

@@ Appendix I Large-scale Experiments
Optimality of the Bounds.
To the best of our knowledge, there are currently no established lower bounds on the convergence rate for the class of gossip-based algorithms applied to either decentralized ranking or trimmed mean estimation. Thus, the optimality of our algorithms remains an open question. In the special case of a complete graph, we observe that significantly faster convergence than the typical $\mathcal{O}(1/t)$ rate is possible. For example, in the Baseline++ algorithm, which performs direct ranking swaps, we can show exponential convergence of the form $\exp(-t/|E|)$. However, this fast convergence critically relies on the high connectivity of the complete graph and does not generalize well to less connected graphs. In contrast, GoRank demonstrates more robust performance across general graphs, including those with limited connectivity, where achieving $\mathcal{O}(1/t)$ convergence is already non-trivial. Deriving a formal lower bound for arbitrary graphs remains an open and challenging problem. Nevertheless, it can be noted that with the current approach, the GoRank convergence rate of $\mathcal{O}(1/t)$ is tight in the sense that rank estimation involves averaging $t$ random indicator functions. Finally, since GoTrim relies on the rank estimates from GoRank, it inherits the $\mathcal{O}(1/t)$ convergence rate, which is already near-optimal.
Faster Gossip Algorithms.
Designing faster gossip algorithms is an interesting direction, and we have also been exploring this aspect. One promising approach involves optimizing the edge sampling strategy based on the graph connectivity constant $c$ (see Xiao and Boyd [2004], Boyd et al. [2004]). While this strategy improves gossip performance in standard averaging tasks, our empirical experiments suggest that its impact on the GoRank algorithm is more limited. This is likely due to the relatively small variation in $c$, which is insufficient to significantly affect a rate in 1/ct1/ct—though it has a more noticeable effect under a geometric rate.
Extension to Multivariate Data.
Extending the proposed approach to multivariate data is an important direction. A natural extension involves ranking multivariate observations based on their norms. Specifically, one could define a suitable norm depending on the task (e.g., Euclidean), compute the norm for each observation, and then apply our univariate ranking method (GoRank) to these one-dimensional values. Observations with the largest norms can then be treated as potential outliers. Following this, a multivariate version of GoTrim can be defined by discarding the top $k=\lfloor\alpha n\rfloor$ observations with the largest norms, and computing the mean of the remaining points using the standard gossip algorithm. Alternatively, one could explore data depths as a generalization of ranking in multivariate settings (see Mosler [2013]). Data depth provides a measure of centrality for multivariate data. While this approach is promising, it would require a more in-depth investigation beyond the scope of the current work, since depth computations usually require to solve computationally demanding optimization problems, just like alternative methods recently designed to define multivariate ranks (e.g. based on optimal transport). Finally, we could simply compute the coordinate-wise trimmed mean by applying GoTrim to each coordinate individually [Pillutla et al., 2022, Yin et al., 2018].
@@
The proposed methods offer a promising foundation for robust decentralized optimization. In particular, GoTrim could be integrated into existing mean-based optimization algorithms to enhance robustness against outliers [Koloskova et al., 2019, Colin et al., 2016, Wu et al., 2018]. Realizing this integration, however, will require further algorithmic and theoretical development. Additionally, our ranking algorithms may prove valuable in extreme value theory, where identifying rare or extreme observations is critical—especially in high-stakes domains such as finance, insurance, and environmental science [Jalalzai et al., 2018].